OpenStudy (anonymous):

The density of a thin circular plate of radius 2 is given by p(x,y)=4+xy.The edge of the plate is described by the parametric equations x=2cost,y=2sint for t in [0,2pi].Find the rate of change of the density with respect to t on the edge of the plate.Does this mean evaluate dx/dt & dy/dt??? or dx/dy which equals (dx/dt)/(dy/dt)?

OpenStudy (blockcolder):

Use the Multivariable Chain Rule and find \[{\partial{p} \over \partial{t}}\]

OpenStudy (anonymous):

sorry for my ignorance.... but i always thought density was a type of rate of change (a derivative). so if you want the rate of change of the density, aren't you actually looking for the second derivative?

OpenStudy (anonymous):

\[dp/dt=dp/dx * dx/dt + dp/dy* dy/dt?\]

OpenStudy (anonymous):

but the equation p(x,y) is defined as the density

OpenStudy (anonymous):

hey you almost spelled my name.. :)

OpenStudy (anonymous):

oh i see...

OpenStudy (anonymous):

thanks.... i didn't read that properly....

OpenStudy (anonymous):

i think you got it... dp/dt... because it does say with respect to t... that would be my guess. sorry can't be of much help.

OpenStudy (blockcolder):

\[\frac{\partial{p}}{\partial{t}}=\frac{\partial{p}}{\partial{x}}\cdot \frac{dx}{dt}+\frac{\partial{p}}{\partial{y}}\cdot\frac{dy}{dt}\]

OpenStudy (anonymous):

is my final answer then \[dp/dt=-4\sin^2t+4\cos^2t?\]

OpenStudy (blockcolder):

Yep. :D

OpenStudy (anonymous):

Thank you blockcolder!!! :-)

OpenStudy (blockcolder):

You're welcome. :D